Pedro has 209 mango trees and star apple trees.If he had 23 more mangoes and 18 fewer star apple trees,the number of mango trees would equal the number of star apple trees. How many mango trees does he have?( answer is 84 show your solution)
Answer by richwmiller(17219)

m+s=209
m+23=s-18
m+s=-23-18
m-s=-41
m+s=209
2m=168
m=84